OpenStudy (anonymous):

Find the future value of an annuity due for which $2,000 payments are made annually for 3 years at 6%.

jimthompson5910 (jim_thompson5910):

use the formula given on this page http://www.financeformulas.net/Future-Value-of-Annuity-Due.html

jimthompson5910 (jim_thompson5910):

Since things are done annually, this means r = 0.06 (6% = 6/100 = 0.06) The annual payments or deposits made are P = 2000 There are n = 3 periods (3 years)